# Testing Monotonicity of Real-Valued Functions on DAGs

### v1

@@ 4.7 A $\Omega(\sqrt{n})$ instantiation and closing mean control
Consider the “shift-matching” scaffold of Section 3.4 built from a set of shifts $A\subseteq\mathbb{Z}^{k}$ and let $U$ be the union graph of the matchings $\{M_{\mathbf{a}}\}_{\mathbf{a}\in A}$. If $A$ is difference-free in the sense that
@@ A $4$-cycle in the underlying undirected graph has the form
The explicit scaffold and trimming.
@@
We extract a balanced dense core by degree-pruning and a simple balancing step. A simultaneously perfect/regular subgraph need not exist for multiple distinct shifts, and this core is sufficient for the barrier parameters.
Lemma 4.24 (Balanced dense core of the $k=2$ shift-matching scaffold).
@@
Now perform the standard degree-pruning procedure: starting from $U$, iteratively delete any vertex (on either side) whose current degree is less than $(c_{0}/4)s$, together with all incident edges, until no such vertex remains. Let the resulting subgraph be $U_{1}=(L_{1},R_{1};E_{1})$. By construction, every vertex in $U_{1}$ has degree at least $(c_{0}/4)s$, and trivially every degree is at most $s$. Each deletion step removes fewer than $(c_{0}/4)s$ edges, so after at most $|L|+|R|=2N^{4}$ steps we have
@@
Consider the parameter choice in Theorem 4.27 (in particular $\alpha=\lambda=1$, $\beta=10$, $B=4\gamma$, $u=\beta\gamma/10$), and assume the scaffold $G$ is $C_{4}$-free with degrees satisfying $\delta\leq\deg(v)\leq\Delta$ for all $v$, where $\Delta/\delta=O(1)$ and $\Delta=\Theta(\sqrt{n})$. Then, for all sufficiently large $n$ (equivalently, all sufficiently large $Q=\Theta(\sqrt{n})$), the mean-control quantity from (18) satisfies
@@
(v) Apply the fixed-budget barrier. Theorem 4.22 states that any deterministic adaptive one-sided tester with query budget $q\leq Q$ must satisfy
@@
Apply Yao’s minimax principle to the hard distribution on instances induced by sampling $I\sim{\rm Unif}[s]$ and then $f\sim\mu_{I}$. Theorem 4.27 gives the corresponding lower bound for deterministic adaptive one-sided testers under this distribution, hence it also bounds randomized testers. ∎

### v2

@@ 4.7 An $\Omega(\sqrt{n})$ instantiation and the mean-control bound
Consider the “shift-matching” base graph of Section 3.4 built from a set of shifts $A\subseteq\mathbb{Z}^{k}$ and let $U$ be the union graph of the matchings $\{M_{\mathbf{a}}\}_{\mathbf{a}\in A}$. If $A$ is difference-free in the sense that
@@ A $4$-cycle in the underlying undirected graph has the form
The explicit base graph and trimming.
@@
We extract a balanced dense subgraph by pruning by degree and a simple balancing step. A simultaneously perfect/regular subgraph need not exist for multiple distinct shifts, and this subgraph is sufficient for the lower-bound parameters.
Lemma 4.24 (Balanced dense subgraph of the $k=2$ shift-matching base graph).
@@
Now perform the standard pruning-by-degree procedure: starting from $U$, iteratively delete any vertex (on either side) whose current degree is less than $(c_{0}/4)s$, together with all incident edges, until no such vertex remains. Let the resulting subgraph be $U_{1}=(L_{1},R_{1};E_{1})$. By construction, every vertex in $U_{1}$ has degree at least $(c_{0}/4)s$, and trivially every degree is at most $s$. Each deletion step removes fewer than $(c_{0}/4)s$ edges, so after at most $|L|+|R|=2N^{4}$ steps we have
@@
Consider the parameter choice in Theorem 4.27 (in particular $\alpha=\lambda=1$, $\beta=10$, $B=4\gamma$, $u=\beta\gamma/10$), and assume the base graph $G$ is $C_{4}$-free with degrees satisfying $\delta\leq\deg(v)\leq\Delta$ for all $v$, where $\Delta/\delta=O(1)$ and $\Delta=\Theta(\sqrt{n})$. Then, for all sufficiently large $n$ (equivalently, all sufficiently large $Q=\Theta(\sqrt{n})$), the mean-control quantity from (18) satisfies
@@
(v) Apply the fixed-budget lower bound. Theorem 4.22 states that any deterministic adaptive one-sided tester with query budget $q\leq Q$ must satisfy
@@
Apply Yao’s minimax principle to the instance distribution induced by sampling $I\sim{\rm Unif}[s]$ and then $f\sim\mu_{I}$. Theorem 4.27 gives the corresponding lower bound for deterministic adaptive one-sided testers under this distribution, hence it also bounds randomized testers. ∎